In a 13 L vessel initially following reaction occur C(s) + S2(g) =CS(g) by 12 g C, 64 g S2, 76 g CSat 1027oC temperature then total pressure is ___.

Grade:9

2 Answers

Sengar Ketan Kumar
6 years ago
Use P=nRT/VSince total pressure at equilibrium is equal to initial partial pressure of S2Ans is 100R.
Preeti
5 years ago
C = 1 mole 
S= 64/32 =2 mole 
CS2 = 76/76=1 mole
C is a solid
Total gaseous mole =2 
PV = nRT 
P= 2× R× 1300/(13)
P=200R